# Proper Name Diacritization for Arabic Wikipedia: A Benchmark Dataset

### v2

@@ Abstract
Proper Name Diacritization for Arabic Wikipedia:
@@ A Benchmark Dataset
Rawan Bondok1, Mayar Nassar1,2, Salam Khalifa3,1, Kurt Micallef4,1, Nizar Habash1 Computational Approaches to Modeling Language (CAMeL) Lab 1New York University Abu Dhabi, 2Ain Shams University 3Stony Brook University, 4Department of Artificial Intelligence, University of Malta {rawan.bondok,nizar.habash}@nyu.edu, mayar.nassar@art.asu.edu.eg salam.khalifa@stonybrook.edu, kurt.micallef@um.edu.mt
@@
Proper Name Diacritization for Arabic Wikipedia:
@@
Rawan Bondok1, Mayar Nassar1,2, Salam Khalifa3,1, Kurt Micallef4,1, Nizar Habash1 Computational Approaches to Modeling Language (CAMeL) Lab 1New York University Abu Dhabi, 2Ain Shams University 3Stony Brook University, 4Department of Artificial Intelligence, University of Malta {rawan.bondok,nizar.habash}@nyu.edu, mayar.nassar@art.asu.edu.eg salam.khalifa@stonybrook.edu, kurt.micallef@um.edu.mt
@@ 1 Introduction
The work presented in this paper lies at the intersection of three commonly but often independently studied Arabic NLP tasks: transliteration, diacritization, and lemmatization. Transliteration focuses on mapping words, primarily proper names, from one script to another, usually in the context of machine translation Beesley (1997); Benites et al. (2020); Chen et al. (2018). It poses challenges due to misalignments between scripts, differences in representing phonological and morphological features, and other historical ad hoc conventions. Diacritization, or diacritic restoration, aims to recover omitted diacritical marks in languages that rely on them for disambiguation Alqahtani et al. (2019); Darwish et al. (2017); Abandah et al. (2015). While both transliteration and diacritization have been well-studied for Arabic, they are typically treated in isolation. One exception is the work of Mubarak et al. (2009), which considers both in the context of Arabic-to-English name transliteration. Lemmatization focuses on mapping inflected words to their base forms. This is particularly important for morphologically rich languages such as Arabic Roth et al. (2008). In the context of Wikipedia entries, providing the lemmas is useful to readers as it gives them grounding on how to interpret and later inflect the word forms properly. More concretely, we focus here on mapping pairs of undiacritized Arabic proper nouns and their English glosses to fully diacritized and lemmatized Arabic forms. The task can be viewed as partial transliteration, where Roman-script vowels help infer (or transliterate into) Arabic diacritical marks. For example, \<نخجوان> nxjwAn ‘Nakhchivan’ (from Figure 1) should ideally be mapped to \<نَخْجِوَان> nax.jiwaAn, rather than incorrect alternatives like \<نِخْجَوَان> nix.jawaAn, or \<نُخْجِوَان> nux.jiwaAn. We present a new dataset of 3,000 unique Arabic Wikipedia proper names annotated with gold lemma-level diacritizations. Each entry is paired with its English Wikipedia equivalent, enabling the study of joint diacritization and transliteration. We benchmark GPT-4o OpenAI et al. (2024), which shows promising results but struggles with spelling variants and ambiguity. The dataset covers a range of named entities (people, places, and organizations) and includes 3,362 total pairs to reflect multiple valid diacritizations based on gloss. Our contributions are:
The remainder of this paper is structured as follows. Section 2 outlines key Arabic linguistic aspects. Section 3 reviews related work. Sections 4 and 5 describe our dataset and annotation process. Section 6 presents evaluation results and error analysis.
@@ 2 Linguistic Background
2.1 Arabic Diacritization
[…]
2.2 Arabic Lemmatization
In Arabic morphology, a lemma is the canonical form (also known as citation form or base word) of a word that abstracts over its inflected variants, including gender, number, person, and case, as well as attached clitics Roth et al. (2008); Habash (2010). Table 2 shows examples of input forms and their corresponding lemmas. In our context, lemmatization is simpler than in free-form text: we focus only on proper nouns, an English gloss is available to guide vowelization, and clitics are rare aside from the definite article (DET) \<ال> Al. The main challenges are distinguishing between baseword and determiner uses of \<ال> Al initial substring (see Table 2 rows 1-2), and handling plural endings (3MP) in demonyms (Table 2 rows 5-6).
@@ 2.3 Arabic Transliterations
Transliteration from Roman script to Arabic presents several challenges, primarily due to the misalignment between the phonological values of the original language and its Roman script orthography, as well as differences between the phonologies of the original languages and Arabic. Arabic, for example, has fewer vowel sounds (6 in Arabic vs. 15 in English), and some missing (no /p/ or /v/) and additional consonants (e.g., emphatic /d/ and /q/). Arabic dialects vary in phonology, including sound quality, letter mapping, and syllabification, leading to multiple valid transliterations. For instance, the borrowed word ‘plastic’ can have different pronunciations and spellings, reflecting variations in vowels and syllabification (see Table 3). During annotation, we adhered to the Wikipedia spelling and aligned closely with the English gloss. The team included Egyptian and Levantine speakers, with the primary annotator being Egyptian.
@@ 3.1 Diacritization in Arabic NLP
Arabic diacritization has been extensively studied using both statistical and neural methods. Some approaches treat it as a standalone task (Zitouni et al., 2006; Mubarak et al., 2019), while others integrate it into multi-task learning frameworks alongside linguistically related tasks such as part-of-speech tagging (Habash and Rambow, 2005; Alqahtani et al., 2020). A commonly adopted strategy involves the use of morphological analyzers. For instance, Camelira implements an analyze-and-disambiguate pipeline: a morphological analyzer generates candidate analyses, which are then ranked by a classifier Obeid et al. (2022). Similarly, Farasa uses morphological patterns to generate the diacritized words Darwish et al. (2017). Systems like Farasa (Darwish et al., 2017) and Camelira Obeid et al. (2022) have demonstrated strong performance sentence-level diacritization tasks. These systems, however, are not directly applicable to our task, which centers on isolated proper nouns, adheres to a task-specific diacritization schema, and incorporates lemma mapping. Unlike sentence-based systems that leverage surrounding context for disambiguation, our task involves context-free diacritization, which poses distinct challenges (see Section 5.1).
@@ 3.2 Lemmatization in Arabic NLP
Lemmatization is another core task in Arabic NLP, and several tools offer robust performance across a variety of syntactic categories (Jarrar et al., 2024; Obeid et al., 2020, 2022). However, our lemmatization task has a narrower scope: it is limited to proper nouns that have a limited inflectional space (see Section 5.1 for further details on our lemmatization space).
@@ 3.3 Transliteration in Arabic NLP
Earlier research on Arabic–English transliteration relied on statistical approaches (Abduljaleel and Larkey, 2004), followed by more targeted work on proper nouns using models such as phonemic memory networks Tian et al. (2022). A persistent challenge in this area is the lack of standardization in transliterating foreign names into Arabic, a problem exacerbated by the omission of diacritics Aziz (1983); Odisho (1992). To address the lack of standardization and limited resources, we introduce a new dataset and annotation guidelines specifically designed for the task of utilizing proper noun transliteration as a signal for Arabic Diacritization. Prior efforts investigated the intersection of transliteration and diacritization, such as Mubarak et al. (2009) and Darwish et al. (2017). Mubarak et al. (2009) used diacritization as a preprocessing step to transliteration. Although, the approach presented in Darwish et al. (2017) for automatically diacritizing transliterated words included leveraging English transliterations to generate Arabic diacritized proper nouns, both their training and test sets were limited in size (500 and 200 instances, respectively). Our resource, by contrast, is publicly available, much larger (3,000 diacritized lemmas), and benchmarked for robust evaluation and development.
@@ 3.4 Arabic Proper Noun Resources
Although various Arabic proper noun datasets exist, they often suffer from limited accessibility, lack of diacritics, or domain constraints. For example, Matthews (2007) compiled a list of 10,001 Arabic names, but the dataset is not publicly available. Eryani and Habash (2021) provide automatically Romanized Arabic bibliographic entries without diacritics, and both the Dan database (Halpern et al., 2009) and SAMA Graff et al. (2009) include diacritized proper nouns, but they were mainly collected form news sources. Khairallah et al. (2024) released a large set of proper nouns as part of their CamelMorph Arabic morphological analyzer (henceforth AraProp). The dataset consists of two distinct portions: (a) AraPropComm, which extends the SAMA Graff et al. (2009) proper-noun list and updates their diacritizations; and (b) AraPropWiki which comprises 63K entries extracted from a Wikidata dump (14-Mar-2023). [^3] https://dumps.wikimedia.org/wikidatawiki/entities/ The AraPropWiki was filtered by Khairallah et al. (2024) to include only single word entities in Arabic and English, and covering only personal and family names, locations and organizations. Unfortunately, Khairallah et al. (2024) did not provide diacritizations for the AraPropWiki portion. Our interest in this topic started by this problem in their open-source resource, which was not usable for our purposes. We discuss these datasets further in Section 4. In this work, we present the first publicly available dataset of maximally manually diacritized and lemmatized Arabic proper nouns on a portion of the AraPropWiki dataset sourced from Wikimedia and manually annotated using English equivalents in a consistent and standardized annotation scheme. To support future work, we also release detailed annotation guidelines and provide the first benchmark of GPT-4’s performance on this task, offering a new resource for evaluating Arabic proper noun diacritization and transliteration.
AraPropComm AraPropWiki AraPropWikiD3K
@@
Table 4: Comparison of dataset statistics across AraPropComm, AraPropWiki, and the annotated subset AraPropWikiD3K.
AraPropComm AraPropWikiD3K
Entries 7,202 3,362
Avg Freq 205,077 61,544
Median Freq 11,732 75
Avg Freeman Similarity Score $\uparrow$ 0.92 $\uparrow$ 0.91
Table 5: Statistics comparing AraPropComm and AraPropWikiD3K subsets. In terms of average frequency (Avg Freq) and median frequency (Med Freq) measured against a frequency list, and Freeman Similarity Score.
@@ 4 Datasets
We work with the AraProp dataset, released as part of an Arabic morphological analyzer by Khairallah et al. (2024). As noted in Section 3, it consists of two parts: AraPropComm and AraPropWiki. We randomly selected 3,000 unique Arabic-script proper nouns from AraPropWiki for manual annotation, forming our dataset AraPropWikiD3K. Table 4 compares the three datasets in terms of unique Arabic entries and full Arabic–English gloss pairs. The original data included multiple glosses per Arabic word (12–20% extra on average). We normalized this by splitting them into separate one-to-one pairs. For example, \<آنا> ĀnA, glossed as ‘A’ana; Ana; Anna’, became three distinct entries: (\<آنا> ĀnA, ‘A’ana’), (\<آنا> ĀnA, ‘Ana’), and (\<آنا> ĀnA, ‘Anna’). Thus, our 3,000 Arabic words expanded to 3,362 Arabic–gloss pairs. Table 5 presents some additional details comparing AraPropComm and AraPropWikiD3K in terms of average and median frequency and Arabic-English phonological similarity. For frequency we used the Arabic Frequency list from Khalifa et al. (2021). For phonological similarity, we used the Freeman similarity score Freeman et al. (2006). While phonological similarity is only slightly lower in AraPropWikiD3K, the overall frequency is significantly lower, highlighting the importance of modeling the diacritization of low-frequency proper nouns in Wikipedia and NLP.
@@
\<سَانْشِيز> sAnšiyz Sanchez Long vowels require preceding diacritics \<سَانْشِيز> saAn.šiyz
@@ 5 Data Annotation
In this section, we discuss the diacritization guidelines we used, as well the setup for initial automatic processing followed by manual correction.
@@ 5.1 Diacritization Guidelines
This effort focuses exclusively on the diacritization of proper nouns and mapping them to their lemmas. As such we require the removal of clitics such as the definite article and the removal of plural suffixes (see Section 2.2).
@@
The final letter \<ي> y has multiple diacritizations that intersect with changes in dialectal Arabic final y-gemination softening (into a /i/). As such, we had to dedicate part of the guidelines to outline the rules for diacritizing it as a geminated /yy/, a long vowel /i:/ or a glide /ay/. The geminated version is the most specific in requirements with three possible cases:
@@ 5.2 Initial Automatic Diacritization
To speed up the annotation process, we gave our annotator an automatically diacirtized version of the data. We used GPT-4o with Arabic Input and English Gloss (comparable to the best setting in Section 6). At the time of creating the initial automatic diacritization, we assumed this a reasonable starting point.
@@
The output of GPT-4o was not always usable as is. When applying well-formedness checks to the diacritized outputs generated by GPT-4o, we observed several recurring patterns of errors that compromised the validity of the diacritized forms. In response, we developed an automated pipeline specifically aimed at correcting these systematic errors. The automatic correction procedures included the following operations:

### v3

# Proper Noun Diacritization for Arabic Wikipedia: A Benchmark Dataset

@@ Abstract
Proper Noun Diacritization for Arabic Wikipedia:
@@ A Benchmark Dataset
Rawan Bondok,1 Mayar Nassar,1,2 Salam Khalifa,1,3 Kurt Micallef,1,4 Nizar Habash1 Computational Approaches to Modeling Language (CAMeL) Lab 1New York University Abu Dhabi, 2Ain Shams University 3Stony Brook University, 4Department of Artificial Intelligence, University of Malta {rawan.bondok,nizar.habash}@nyu.edu, mayar.nassar@art.asu.edu.eg salam.khalifa@stonybrook.edu, kurt.micallef@um.edu.mt
@@
Proper Noun Diacritization for Arabic Wikipedia:
@@
Rawan Bondok,1 Mayar Nassar,1,2 Salam Khalifa,1,3 Kurt Micallef,1,4 Nizar Habash1 Computational Approaches to Modeling Language (CAMeL) Lab 1New York University Abu Dhabi, 2Ain Shams University 3Stony Brook University, 4Department of Artificial Intelligence, University of Malta {rawan.bondok,nizar.habash}@nyu.edu, mayar.nassar@art.asu.edu.eg salam.khalifa@stonybrook.edu, kurt.micallef@um.edu.mt
@@ 1 Introduction
The work presented in this paper lies at the intersection of three commonly but often independently studied Arabic NLP tasks: transliteration, diacritization, and lemmatization. Transliteration is the mapping of words, primarily proper nouns, from one script to another, usually in the context of machine translation Beesley (1997); Benites et al. (2020); Chen et al. (2018). It poses challenges due to misalignments between scripts, differences in representing phonology and morphology, and historical ad hoc conventions. Diacritization, or diacritic restoration, aims at recovering omitted diacritics in languages that rely on them for disambiguation Alqahtani et al. (2019); Darwish et al. (2017); Abandah et al. (2015). While both transliteration and diacritization have been well studied for Arabic, they are typically treated in isolation. An exception is the work of Mubarak et al. (2009), which considers both in the context of Arabic to English proper noun transliteration. Lemmatization maps inflected words to their base forms. This is particularly important for morphologically rich languages such as Arabic Roth et al. (2008). In the context of Wikipedia entries, providing the lemmas is useful to readers as it gives them a grounding on how to interpret and later inflect the word forms properly. More concretely, we focus here on mapping pairs of undiacritized Arabic proper nouns and their English glosses to fully diacritized and lemmatized Arabic forms. The task can be viewed as partial transliteration, where Roman-script vowels help infer (or transliterate into) Arabic diacritical marks. For example, \<نخجوان> nxjwAn ‘Nakhchivan’ (from Figure 1) should ideally be mapped to \<نَخْجِوَان> nax.jiwaAn, rather than incorrect alternatives like \<نِخْجَوَان> nix.jawaAn or \<نُخْجِوَان> nux.jiwaAn. We present a new dataset of 3,000 unique Arabic Wikipedia proper nouns annotated with gold lemma-level diacritizations. Each entry is paired with its English Wikipedia equivalent, enabling the study of joint diacritization and transliteration. We benchmark GPT-4o OpenAI et al. (2024), which shows promising results but struggles with spelling variants and ambiguity. The dataset covers a range of named entities (people, places, and organizations) and includes 3,362 total pairs to reflect multiple valid diacritizations based on the gloss. Our contributions are:
  • A publicly available gold-standard dataset of Arabic Wikipedia proper nouns with English equivalents.LABEL:footCP
  • A GPT-4o benchmark and detailed error analysis for Arabic proper noun diacritization.
The remainder of this paper is structured as follows. Section 2 outlines Arabic linguistic aspects. Section 3 reviews related work. Sections 4 and 5 describe our dataset and annotation process. Section 6 presents evaluation results and error analysis.
@@ 2 Linguistic Background
2.1 Arabic Diacritization
[…]
2.2 Arabic Lemmatization
In Arabic morphology, the lemma is the canonical form (also known as citation form) of a word that abstracts over its inflected variants, including gender, number, person, and case, as well as attached clitics Roth et al. (2008); Habash (2010). Table 2 shows examples of input forms and their corresponding lemmas. In our context, lemmatization is simpler than in free-form text: we focus only on proper nouns, an English gloss is available to guide vowelization, and clitics are rare. The main challenges are distinguishing between baseword and determiner uses of \<ال> Al (DET) initial substring (see Table 2 rows 1-2), and handling plural endings (3MP) \<ون> uw in demonyms (Table 2 rows 5-6).
@@ 2.3 Arabic Transliterations
Transliteration from Roman script to Arabic script presents several challenges, primarily due to the misalignment between the phonology of the original language and its Roman script orthography, as well as differences between the phonology of the original languages and Arabic. Arabic, for example, has fewer vowels (6 in Arabic vs. 15 in English), and some missing (no /p/ or /v/) and additional consonants (e.g., emphatic /d/ and /q/). Arabic dialects vary in phonology, including sound quality, letter mapping, and syllabification, leading to multiple valid transliterations. For instance, the borrowed word ‘plastic’ can have different pronunciations and spellings, reflecting variations in vowels and syllabification (see Table 3). During annotation, we followed Wikipedia spelling and aligned with the English gloss. The team included Egyptian, Sudanese, and Levantine speakers, with an Egyptian speaker as the primary annotator.
@@ 3.1 Diacritization in Arabic NLP
Arabic diacritization has been extensively studied using both statistical and neural methods. Some approaches treat it as a standalone task (Zitouni et al., 2006; Mubarak et al., 2019), while others integrate it into multitask learning frameworks alongside linguistically related tasks such as part-of-speech tagging (Habash and Rambow, 2005; Alqahtani et al., 2020). A commonly adopted strategy involves the use of morphological analyzers. For instance, Camelira (Camel Tools) implements an analyze-and-disambiguate pipeline: a morphological analyzer generates candidate analyses, which are then ranked by a classifier Obeid et al. (2020, 2022). Similarly, Farasa uses morphological patterns to diacritize words Darwish et al. (2017). Systems such as Farasa and Camel Tools have demonstrated strong performance on sentence-level diacritization tasks. However, these systems are not directly applicable to our task, which centers on isolated proper nouns, adheres to a task-specific diacritization schema, and incorporates lemma mapping. Unlike sentence-based systems that leverage surrounding context for disambiguation, our task involves context-free diacritization, which poses distinct challenges (see Section 5.1).
@@ 3.2 Lemmatization in Arabic NLP
Lemmatization is another core task in Arabic NLP, and several tools offer robust performance across a variety of syntactic categories (Obeid et al., 2020, 2022; Jarrar et al., 2024). However, our lemmatization task has a narrower scope: it is limited to proper nouns that have a limited inflectional space (see Section 5.1 for further details on our lemmatization space).
@@ 3.3 Transliteration in Arabic NLP
Earlier research on Arabic–English transliteration relied on statistical approaches (Abduljaleel and Larkey, 2004), followed by more targeted work on proper nouns using models such as phonemic memory networks Tian et al. (2022). A persistent challenge in this area is the lack of standardization in transliterating foreign names into Arabic, a problem exacerbated by the omission of diacritics Aziz (1983); Odisho (1992). To address the lack of standardization and limited resources, we introduce a new dataset and annotation guidelines specifically designed for the task of utilizing proper noun transliteration as a signal for Arabic diacritization. Prior efforts investigated the intersection of transliteration and diacritization, such as Mubarak et al. (2009) and Darwish et al. (2017). Mubarak et al. (2009) used diacritization as a preprocessing step to transliteration. Although, the approach presented in Darwish et al. (2017) for automatically diacritizing transliterated words included leveraging English transliterations to generate Arabic diacritized proper nouns, both their training and test sets were limited in size (500 and 200 instances, respectively). Our resource, in contrast, is publicly available, much larger (3,000 diacritized lemmas), and benchmarked for robust evaluation and development.
@@ 3.4 Arabic Proper Noun Resources
Although various Arabic proper noun datasets exist, they often suffer from limited accessibility, lack of diacritics, or domain constraints. For example, Matthews (2007) compiled a list of 10,001 Arabic names, but the dataset is not publicly available. Eryani and Habash (2021) provide automatically Romanized Arabic bibliographic entries without diacritics, and both the Dan database (Halpern et al., 2009) and SAMA Graff et al. (2009) include diacritized proper nouns, but they were mainly collected form news sources. Khairallah et al. (2024) released a large set of proper nouns as part of their CamelMorph Arabic morphological analyzer (henceforth CamelProp, CP for short). The dataset consists of two distinct portions: (a) CP-Sama, which extends the SAMA Graff et al. (2009) proper-noun list and updates their diacritizations; and (b) CP-Wiki which comprises 63K entries extracted from a Wikidata dump (14-Mar-2023). [^3] https://dumps.wikimedia.org/wikidatawiki/entities/ The CP-Wiki was filtered by Khairallah et al. (2024) to include only single word entities in Arabic and English, and covering only personal and family names, locations and organizations. Unfortunately, Khairallah et al. (2024) did not provide diacritizations for the CP-Wiki portion. Our interest in this topic started by this problem in their open-source resource, which was not usable for our purposes. We discuss these datasets further in Section 4. In this work, we present the first publicly available dataset of maximally manually diacritized and lemmatized Arabic proper nouns on a portion of the CP-Wiki dataset sourced from Wikimedia and manually annotated using English equivalents in a consistent and standardized annotation scheme. To support future work, we also release detailed annotation guidelines and provide the first benchmark of GPT-4o’s performance on this task, offering a new resource for evaluating Arabic proper noun diacritization and transliteration.
CP-Sama CP-Wiki CP-Wiki-D3K
@@
Average Freq 205,077 97,438 61,544
Median Freq 11,732 87 75
Average Freeman Score 0.92 0.91 0.91
@@
Table 4: Comparison of dataset statistics across CP-Sama, CP-Wiki, and the annotated subset CP-Wiki-D3K.
Class CP-Wiki CP-Wiki-D3K
Location 77.1% 85.2%
Name 25.5% 35.0%
Organization 2.0% 2.0%
Table 5: Distribution of different named entity classes across CP-Wiki and CP-Wiki-D3K
@@ 4 Datasets
We work with the CamelProp dataset, released as part of CamelMorph, an Arabic morphological analyzer, by Khairallah et al. (2024). As noted in Section 3, it consists of two parts: CP-Sama and CP-Wiki. We randomly selected 3,000 unique Arabic-script proper nouns from CP-Wiki for manual annotation, forming our dataset CP-Wiki-D3K. Table 4 compares the three datasets in terms of unique Arabic entries and full Arabic–English gloss pairs, average and median frequency and Arabic-English phonological similarity. For frequency we used the Arabic Frequency list from Khalifa et al. (2021). For phonological similarity, we used the Freeman similarity score Freeman et al. (2006). The original data included multiple glosses per Arabic word (12–20% extra on average). We normalized this by splitting them into separate one-to-one pairs. For example, \<آنا> ĀnA, glossed as ‘A’ana; Ana; Anna’, became three distinct entries: (\<آنا> ĀnA, ‘A’ana’), (\<آنا> ĀnA, ‘Ana’), and (\<آنا> ĀnA, ‘Anna’). Thus, our 3,000 Arabic words expanded to 3,362 Arabic–gloss pairs. While phonological similarity is only slightly lower in CP-Wiki-D3K and CP-Wiki, the overall frequency in CP-Wiki and CP-Wiki-D3K is significantly lower than CP-Sama, highlighting the importance of modeling the diacritization of low-frequency proper nouns in Wikipedia and NLP. In addition to frequency and phonological similarity, we examined the distribution of named entity categories, namely, personal and family names, locations, and organizations, across both the original CP-Wiki dataset and the manually annotated subset, CP-Wiki-D3K. The distributions were broadly similar, with location entities being the majority in both (CP-Wiki: 77.1%, CP-Wiki-D3K: 85.2%), followed by names and organizations. This consistency supports the representativeness of CP-Wiki-D3K for studying diacritization across entity types. Table 5 reports the detailed percentage breakdown of entity classes in both datasets.
@@
\<سانشِيز> sAnšiyz Sanchez Long vowels require preceding diacritics \<سَانْشِيز> saAn.šiyz
@@ 5 Data Annotation
In this section, we discuss the diacritization guidelines we used, as well as the setup for initial automatic processing followed by manual correction.
@@ 5.1 Diacritization Guidelines
This effort focuses exclusively on the diacritization of proper nouns and mapping them to their lemmas. As such, we require the removal of clitics such as the definite article and the removal of plural suffixes (see Section 2.2).
@@
The final letter \<ي> y has multiple diacritizations that overlap with changes in dialectal Arabic, i.e. the softening of final y-gemination into /i/. As such, we had to dedicate part of the guidelines to outline the rules for diacritizing it as a geminated /yy/, a long vowel /i:/ or a glide /ay/. The geminated version is the most specific in requirements with three possible cases:
@@ 5.2 Initial Automatic Diacritization
To speed up the annotation process, we gave our annotator an automatically diacirtized version of the data. We used GPT-4o with Arabic Input and English Gloss (comparable to the best setting in Section 6). At the time of generating the initial automatic diacritization, we considered this a reasonable starting point.
Type of Disagreement Freq Gloss First Annotator Second Annotator
Kasra $\leftrightarrow$ Sukun 13 Tibet \<تِبِت> tibit \<تِبْت> tib.t
Kasra $\leftrightarrow$ Fatha 9 Shechem \<شِكِيم> šikiym \<شَكِيم> šakiym
Consonant $\leftrightarrow$ Long vowel 9 Jane \<جَيْن> jay.n \<جِين> jiyn
Sukun $\leftrightarrow$ Damma 5 Acquaviva \<أَكُوَافِيفَا> ÂakuwaAfiyfaA \<أَكْوَافِيفَا> Âak.waAfiyfaA
Sukun $\leftrightarrow$ Fatha 1 Aminadav \<عَمِينْدَاف> $\varsigma$amiyn.daAf \<عَمِينَدَاف> $\varsigma$amiynadaAf
Shadda $\leftrightarrow$ $\phi$ 1 Oss \<أُوسّ> Â$\sim$ \<أُوس> Âuws
Table 7: Types of disagreements in Inter-Annotator Evaluation
@@
The output of GPT-4o was not always usable as is. When applying well-formedness checks to the diacritized outputs generated by GPT-4o, we observed several recurring patterns of errors that compromised the validity of the diacritized forms. In response, we developed an automated pipeline specifically aimed at correcting these systematic errors.LABEL:footCP The automatic correction procedures included the following operations: